OpenStudy (anonymous):

Solve the absolute value equation |2n + 15| = 1. One solution is n = -7. What is the other solution

myininaya (myininaya):

Solve this -(2n+15)=1 2n+15=-1 2n=-1-15 2n=-16 n=-8